OpenStudy (anonymous):

I have to prove this is in an identity. I dont think this one is so if someone could double check for me I would really appreciate it!! cos3t= cos^3(t)sin(t)- sin^3(t)cos(t) does the cubes make this a doable problem?

OpenStudy (perl):

start out with cos(3t) = cos( 2t + t )

OpenStudy (anonymous):

cos 3t=cos(t+2t) now expand this using the identity cos(a+b)=cos a cos b-sin a sin b

OpenStudy (anonymous):

@theee_queenbee

OpenStudy (anonymous):

sorry I am having connection issues :( but when I tried it on paper I got cos(t)sin^2(t)-sin(t)cos^2(t)

OpenStudy (anonymous):

so what am I doing wrong?

OpenStudy (anonymous):

@Princer_Jones

OpenStudy (anonymous):

@theee_queenbee are you proceeding in this manner?

OpenStudy (anonymous):

no I was looking at another identity. thats why I was confused. ok so now I have cos(t) (2cos^2(t)-1) - sin(t) (2sin(t)cos(t)) ( im visual so I write it out the long way unfortunately)

OpenStudy (anonymous):

what should we prove??

OpenStudy (anonymous):

Do you know this how to prove?? cos 3x= 4 cos^3 (x) - 3cos x @theee_queenbee